• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ransome

    Log entry time 00:08:23 on June 4,2002

    Entry number 85189

    keyword=shift summary

    SL - Ransome 16-20, Reitz 20-24
    16:00 begin shift with no stable beam - beam is on and off
    we feel it is not stable enough to use for data taking
    16:20 MCC calls - says they won't be able to get above 45 uA,
    want to take 10 more minutes for some optics check
    16:30 MCC says they have IOC problem will stop beam in
    a few minutes for 15 minutes
    16:45 PD says experts will come at 17:30 for optics checks,
    which may take "a few hours"
    17:20 MCC calls - experts are here, will try to get the show
    back on the road.
    18:30 MCC says they will try to deliver beam "in a few minutes"
    18:33 some beam
    18:45 running with 96 uA
    18:48 beam off
    18:49 beam back
    18:49:30 beam off
    18:51 beam back
    19:30 harp scan, run # 1767, coda run #1132
    20:00 start Q1_D40
    20:25 target IOC reboot (in the middle of 1136)
    21:30 lots of beamtrips, some runs are junk
    23:50 change to Q1D_d50